# arxiv-appendix

### v1

@@ 2.3.1 Overview
We start with a skeleton of our proposed method. Due to the nonparametric nature of the problem, we dynamically partition the covariate space $\mathcal{X}$ into a set of hypercubes (bins) and employ a locally constant estimator, subject to LDP, of the reward functions. The partition strategy converts the contextual problem into a collection of static MAB decision problems, which are then dealt with using a confidence bound-type arm-elimination method. Given that all arms are pulled sufficiently, we can identify and eliminate those sub-optimal arms based on local estimates and confidence bounds.
We first introduce the policy $\pi_{t}$ used at time step $t\in[n_{\mathrm{P}}]$. Specifically, we maintain an active partition $\mathcal{B}_{t}$, initialized as $\mathcal{B}_{0}=\{B_{0}^{0}\}$ and updated dynamically. The subscript and superscript of $B_{s}^{j}$ denote the depth and index, respectively, which will be explained in detail later. For each $B_{s}^{j}$, define $A_{s}^{j}\subset[K]$ as the candidate arms set. Upon observing a new covariate $X_{t}^{\mathrm{P}}$, belonging to some $B_{s}^{j}\in\mathcal{B}_{t}$, we select
$\pi_{t}(x)=\operatorname{Unif}(A_{s}^{j}),$ (8)
@@
We now elaborate on the policy updating, consisting of three main steps and illustrated in Figure 2.
@@
    Update the private local estimates for reward functions. As shown in Figure 2(a), in each bin, there are $|A_{s}^{j}|$ active arms, each with its own estimate. We design a mechanism to optimally estimate the reward functions under the LDP constraint. This step is formulated in Section 2.3.3.
@@
    Decide if any arm needs to be eliminated. Using a confidence bound for LDP nonparametric contextual MAB, we identify and remove suboptimal arms. This step is illustrated in Figure 2(b) and formulated in Section 2.3.4.
@@
    Decide whether the current bin should be refined. If some arm $k$ is pulled too many times inside $B_{s}^{j}$, the confidence bound becomes narrow. In this case, the ability to distinguish sub-optimal arms is restricted by the approximation error of $B_{s}^{j}$, which can then be improved by refining the bin to sub-bins. This step is illustrated in Figure 2(c) and formulated in Section 2.3.5.
@@
The complete procedure is detailed in Algorithm 2. For the clarity of presentation, before discussing the three key steps, we first introduce the dynamic partitioning, i.e. the placement of the dashed lines in Figure 2(c), in Section 2.3.2.
@@ 2.3.2 Dynamic Partitioning
A partition of domain $\mathcal{X}$ is a collection of nonempty, pairwise disjoint subsets whose union is $\mathcal{X}$. To create a partitioning of $\mathcal{X}$, let the rectangular bin at the root level be $B_{0}^{0}:=[0,1]^{d}$. For each bin $B_{s}^{j}$, where $s$ represents its depth and $j\in[2^{s}]$ is its index, two successive bins are created in the following way. In particular, we uniformly choose a dimension among those embedding longest edges of $B_{s}^{j}$, then split $B_{s}^{j}$ along this dimension at the midpoint, resulting in sub-bins $B_{s+1}^{2j-1}$ and $B_{s+1}^{2j}$. The partition process is illustrated in Figure 3 and formalized in Algorithm 1. This procedure is widely used in the literature and is referred to as dyadic partition or max-edge partition (e.g. Blanchard et al., 2007; Cai et al., 2024; Ma et al., 2025).
@@
$\mathcal{M}_{sj}=\mathop{\mathrm{argmax}}_{k}|b_{sj}^{k}-a_{sj}^{k}|$.
$k^{*}=$ Unif $(\mathcal{M}_{sj})$.
$B_{s+1}^{2j-1}=\{x|x\in B_{s}^{j},x^{k^{*}}<({a_{sj}^{k^{*}}+b_{sj}^{k^{*}}})/{2}\}$.
$B_{s+1}^{2j}=B_{s}^{j}/B_{s+1}^{2j-1}$.
Output: Sub-bin $B_{s+1}^{2j-1},B_{s+1}^{2j}$.
@@
Figure 3: A partitioning created by the max-edge rule for $d=2$. Blue areas represent the corresponding bins.
@@ 2.3.3 Estimating Reward Functions
In this section, we study the estimation of reward functions subject to LDP constraints. Specifically, we focus on partition-based LDP estimators that assign a constant prediction value within each partition bin. We begin by investigating the non-private counterpart of the partition-based estimation. It simply averages the rewards of data points whose covariates fall into the same bin. We then inject the LDP ingredient and present the final estimator.
Let $a_{t,s}^{j}=1$ if $B_{s}^{j}$ is in $\mathcal{B}_{t}$ and 00 otherwise, with $a_{t,s}^{j}$ being the indicator of whether the bin $B_{s}^{j}$ is in the active partition $\mathcal{B}_{t}$ at time $t$. The size of the sample used to update bin $B_{s}^{j}$ is therefore
$t_{s}^{j}=\sum_{i=1}^{t}a_{i,s}^{j}.$ (9)
Note that both $t_{s}^{j}$ and $a_{i,s}^{j}$ are free of privacy concerns since the server is aware of the active partition (indicated by $\pi_{t}$) at each time step. See the illustration in Figure 1, where $\pi_{t}$ at each step is publicly available. In this case, a non-private estimator for $f^{\mathrm{P}}_{k}$ in $B_{s}^{j}$ is
@@
which is the average of labels in each bin that come from arm $k$. From here on, we define $/0=0$ 0.
To privatize the data, we estimate the reward function under LDP using a Laplace mechanism (Dwork et al., 2006), which is also considered in Berrett and Butucea (2019), Berrett and Yu (2021), Györfi and Kroll (2023), Ma and Yang (2024), among many others. Specifically, there are three components in (10), namely $Y_{i}^{\mathrm{P},(\pi_{i}(X_{i}^{\mathrm{P}}))}$, $\boldsymbol{1}(X_{i}^{\mathrm{P}}\in B^{j}_{s})$ and $\boldsymbol{1}(\pi_{i}(X_{i}^{\mathrm{P}})=k)$, that require privatization. We denote the non-private information by
@@
for $B_{s}^{j}\in\mathcal{B}_{i}$ and $k\in[K]$. Consequently, they are privatized as
@@
where $\xi$’s and $\zeta$’s are i.i.d. standard Laplace random variables. The privacy budget $\varepsilon$ is divided into two parts for privacy preservation on $V$’s and $U$’s, respectively. Note that all $B_{s}^{j}\in\mathcal{B}_{i}$ receives an update based on $Z_{i}^{\mathrm{P}}$ regardless whether $X_{i}^{\mathrm{P}}\in B_{s}^{j}$ or not. Otherwise, bin $B_{s}^{j}$ not receiving an update reveals $X_{i}^{\mathrm{P}}\notin B_{s}^{j}$, which is a privacy leakage. The final estimator is therefore
@@
which satisfies the $\varepsilon$-LDP constraint, as demonstrated in Proposition LABEL:prop:privacyguarantee in appendix.
@@ 2.3.4 Eliminating Arms
Policy (8) uniformly pulls active arms, whereas an ideal policy should exclude arms with large regret from $A_{s}^{j}$. To achieve this, we dynamically rule out arms that are deemed suboptimal in each bin. By a suboptimal arm in a bin, we mean an arm whose reward function is lower than that of another arm for all $x$ in the bin. Although this is an unobservable population property, it can be inferred using a sufficient condition provided in the following proposition. This proposition establishes a bound between the private estimator (13) and its population counterpart $\mathbb{E}_{Y|X,\pi}\left[\widehat{f}^{\mathrm{P},t}_{k}(x)\right]$, defined as
@@
The bound will guide the choice of confidence bound in our method.
@@
For $t\in[n_{\mathrm{P}}]$, let $\mathcal{B}_{t}$ be the output of Algorithm 1. Let $t_{s}^{j}=\sum_{i=1}^{t}a_{i,s}^{j}$ be defined in (9). Then, with probability at least $1-n_{\mathrm{P}}^{-2}$, it holds that, with $t_{s}^{j}\geq\log^{2}(n_{\mathrm{P}})$ and $C_{n_{\mathrm{P}}}\asymp\log(n_{\mathrm{P}})$,
@@
for all $k\in A_{s}^{j},t\in[n_{\mathrm{P}}],B_{s}^{j}\in\mathcal{B}_{t}$, and $x\in B^{j}_{s}\in\mathcal{B}_{t}$.
The proof of Proposition 2.7 can be found in Appendix LABEL:sec:error-analysis, where we also specify the exact expression of $C_{n_{\mathrm{P}}}$. We remark that this proposition is the very first confidence bound result for LDP nonparametric contextual MAB. In the numerator of (14), the two terms correspond to the private and non-private bounds, respectively. The private term $\varepsilon^{-2}t_{s}^{j}$ arises from the sum of Laplace random variables. As for the non-private term, a more natural form is $\sum_{i=1}^{t}{U}_{i,k,s}^{\mathrm{P},j}a^{j}_{i,s}$, which, however, is not accessible. Since our algorithm requires the realization of $r_{k,s}^{\mathrm{P},t,j}$, we replace this bound with its private counterpart $\sum_{i=1}^{t}\tilde{U}_{i,k,s}^{\mathrm{P},j}a^{j}_{i,s}$. Unlike standard confidence bounds, which count the number of samples whose covariates lie in a particular bin during a given time period, our construction introduces additional randomness to comply with LDP requirements. In Lemma LABEL:lem:orderofsumU in the appendix, we theoretically show that this substitution does not compromise the effectiveness of the confidence bound, provided that $t_{s}^{j}$ is sufficiently large. To ensure this condition is met, we design a sufficient exploration criterion when conducting the elimination, which is introduced below.
We derive the elimination rule from (14). By the triangle inequality, it holds $|\tilde{f}_{k}^{\mathrm{P},t}(x)-f_{k}(x)|\leq 2r_{k,s}^{\mathrm{P},t,j}$ provided that
$|\mathbb{E}_{Y|X}\left[\widehat{f}^{\mathrm{P}}_{k}(x)\right]-f_{k}(x)|\leq r_{k,s}^{\mathrm{P},t,j}.$ (15)
Condition (15) is ensured by designing a bin-refining process, which will be discussed in the next section. Then, we can set $2r_{k,s}^{\mathrm{P},t,j}$ as the radius of confidence bound of $\tilde{f}_{k}^{\mathrm{P},t}(x)$. Following conventional confidence bound-based methods, we eliminate an arm when its upper confidence bound is smaller than the lower confidence bound of another arm. Formally, we remove arm $k^{*}$ from $A_{s}^{j}$ if there exists $k\in[K]$ such that
@@
where the first condition ensures that the bin is sufficiently explored. Here, the sufficiently explored condition is required by Proposition 2.7.
@@ 2.3.5 Refining Bins
We describe how to refine the current bin to ensure the claimed condition (15). As explained in Section 2.3.2, this condition implies that the ability to distinguish sub-optimal arms is not dominated by the approximation error of $B_{s}^{j}$. We choose
@@
as a surrogate for the approximation error. Note that the approximation error is decreasing with $s$, i.e. the finer bins have smaller errors. In fact, (17) is the diameter of $B_{s}^{j}$ and represents an upper bound on the approximation error up to a constant factor, as shown in Lemma LABEL:lem:approximationerror in the appendix. When $r_{k,s}^{t,j}<\tau_{s}$, it signals insufficient approximation capability, prompting the refinement of the bin using Algorithm 1.
Input: Budget $\varepsilon$.
Initialization: $\pi_{0}=\operatorname{Unif}([K])$, $\mathcal{B}_{0}=[0,1]^{d}$, $A_{0}^{0}=[K]$, $t=0$.
while True do
@@
             Get $\tilde{V}_{t,k,s}^{\mathrm{P},j}$ and $\tilde{U}_{t,k,s}^{\mathrm{P},j}$.
             Compute estimates (13) and confidence bounds (14). # estimating reward functions
             Remove $k$ from $A_{s}^{j}$ if (16). # eliminating arms
             if  $r_{k,s}^{t,j}<{\tau}_{s}$ for any $k$  then
@@
                   $\mathcal{B}_{t}=\mathcal{B}_{t}\cup\{B_{s+1}^{2j-1},B_{s+1}^{2j}\}/B_{s}^{j}$. # refining bins
@@
      $t\leftarrow t+1$. $\mathcal{B}_{t}=\mathcal{B}_{t-1}$. Update $\pi_{t}$ by (8). Send $\pi_{t}$ to user $t$.
end while
Algorithm 2 LDP MAB

### v2

@@ 2.3.1 Overview
To start, we first provide an overview of our proposed method (see the detailed procedure in Algorithm 2 later) in this subsection. Due to the nonparametric nature of the problem, we dynamically partition the covariate space $\mathcal{X}$ into a set of hypercubes (i.e. bins) and employ a locally constant estimator, subject to LDP, of the reward functions. The partition strategy converts the contextual problem into a collection of static MAB decision problems, which are then dealt with via a confidence bound based arm elimination procedure. In particular, given that all arms are pulled sufficiently, we can identify and eliminate sub-optimal arms based on local estimates and the corresponding confidence bounds. Furthermore, to ensure the approximation error due to binning is negligible, the partition is dynamically updated via a refinement procedure.
The main structure of our algorithm is inspired by the adaptive binning and successive elimination (ABSE) procedure proposed in Perchet and Rigollet, (2013) for non-private nonparametric contextual MAB. However, to accommodate the LDP constraints, substantial modifications are needed on the design of the mechanism for user-server information separation, and on the construction of the nonparametric reward function estimator and its confidence bound. We refer to Section 2.3.6 for a more detailed comparison with Perchet and Rigollet, (2013).
To proceed, we introduce the policy $\pi_{t}$ used at time $t\in[n_{\mathrm{P}}]$. Specifically, we maintain an active partition $\mathcal{B}_{t}$, initialized as $\mathcal{B}_{1}=\{B_{0}^{1}:=[0,1]^{d}\}$ (i.e. the entire covariate domain) and updated dynamically. The subscript and superscript of $B_{s}^{j}$ denote the depth and index of the bin, respectively, which will be explained in detail later. For each bin $B_{s}^{j}$, denote $A_{s}^{j}\subseteq[K]$ as its active arms set. Upon observing a new covariate $X_{t}^{\mathrm{P}}$, belonging to some $B_{s}^{j}\in\mathcal{B}_{t}$, the policy prescribes
$\pi_{t}(X_{t}^{P})=\operatorname{Unif}(A_{s}^{j}),$ (8)
@@
We now elaborate on how the policy $\pi_{t}$ is updated across time, which consists of three key components and is further illustrated in Figure 2. The detailed procedure is given in Algorithm 2.
@@
    Update the private local estimates of the reward functions. As shown in Figure 2(a), in each bin, there are $|A_{s}^{j}|$ active arms, each with its own estimate. We design a mechanism to optimally estimate the reward functions under LDP. This step is formulated in Section 2.3.3.
@@
    Decide if any arm needs to be eliminated. Via a novelly constructed confidence bound for LDP nonparametric contextual MAB, we identify and remove suboptimal arms for each bin in the active partition. This step is illustrated in Figure 2(b) and formulated in Section 2.3.4.
@@
    Decide whether a given bin should be refined. For any active arm in a given bin, the confidence bound for the local estimate of its reward function becomes narrower as the arm is pulled more times. When the confidence bound is sufficiently narrow, the ability to distinguish sub-optimal arms is restricted by the approximation error of the bin, which can then be improved by refining it to sub-bins. This step is illustrated in Figure 2(c) and formulated in Section 2.3.5.
@@
For clarity of presentation, before discussing the three key components, we first detail the partitioning procedure itself, i.e. the placement of the dashed lines in Figure 2(c), in Section 2.3.2.
@@ 2.3.2 Dynamic Partitioning
A partition of domain $\mathcal{X}$ is a collection of nonempty, pairwise disjoint subsets whose union is $\mathcal{X}$. To create a partition of $\mathcal{X}$, let the rectangular bin at the root level be $B_{0}^{0}:=[0,1]^{d}$. For each bin $B_{s}^{j}$, where $s$ represents its depth and $j\in[2^{s}]$ is its index, two successive sub-bins are created in the following way. In particular, we uniformly choose a dimension among those embedding longest edges of $B_{s}^{j}$, then split $B_{s}^{j}$ along this dimension at the midpoint, resulting in sub-bins $B_{s+1}^{2j-1}$ and $B_{s+1}^{2j}$. The partition process is illustrated in Figure 3 and formalized in Algorithm 1. This procedure is widely used in the literature and is referred to as dyadic partition or max-edge partition (e.g. Blanchard et al.,, 2007; Cai et al.,, 2024; Ma et al.,, 2025).
@@
1. Collect $\mathcal{M}_{sj}=\mathop{\mathrm{argmax}}_{k}|b_{sj}^{k}-a_{sj}^{k}|$ and set $k^{*}=$ Unif $(\mathcal{M}_{sj})$.
2. Set $B_{s+1}^{2j-1}=\Big{\{}x:x\in B_{s}^{j},x^{k^{*}}<({a_{sj}^{k^{*}}+b_{sj}^{k^{*}}})/{2}\Big{\}}$ and $B_{s+1}^{2j}=B_{s}^{j}/B_{s+1}^{2j-1}$.
Output: Sub-bins $B_{s+1}^{2j-1},B_{s+1}^{2j}$.
@@
Figure 3: A partition created by the max-edge rule for $d=2$. Blue areas give the corresponding bins.
@@ 2.3.3 Estimating Reward Functions
In this section, we study the estimation of reward functions subject to LDP constraints. Specifically, we focus on partition-based LDP estimators that assign a constant within each partition bin. To build intuition, we begin by investigating the non-private counterpart of this partition-based estimation. It simply averages the rewards of data points whose covariates fall into the same bin. We then inject the LDP ingredient and present the final estimator.
Let $a_{t,s}^{j}=1$ if $B_{s}^{j}$ is in $\mathcal{B}_{t}$ and 00 otherwise. In other words, $a_{t,s}^{j}$ is the indicator of whether the bin $B_{s}^{j}$ is in the active partition $\mathcal{B}_{t}$ at time $t$. We further define
$t_{s}^{j}=\sum_{i=1}^{t}a_{i,s}^{j},$ (9)
which records the total number of times that $B_{s}^{j}$ is in the active partition up to time $t.$ Note that both $t_{s}^{j}$ and $a_{i,s}^{j}$ are free of privacy concerns since the server is aware of the active partition $\mathcal{B}_{t}$ at each time step. Recall the illustration in Figure 1, where $\pi_{t}$ (and thus its associated active partition $\mathcal{B}_{t}$) at each step is publicly available. In this case, a non-private estimator for $f_{k}:=f^{\mathrm{P}}_{k}$ at time $t$ is
@@
which is simply the sample average of all rewards (i.e. $Y$) that come from arm $k$ with their covariates falling into the same bin in $\mathcal{B}_{t}$. Henceforth, we define $/0=0$ 0.
For privacy protection, we estimate the reward function under LDP via the Laplace mechanism (Dwork et al.,, 2006). Specifically, there are three components in (10), namely $Y_{i}^{\mathrm{P},(\pi_{i}(X_{i}^{\mathrm{P}}))}$, $\boldsymbol{1}(X_{i}^{\mathrm{P}}\in B^{j}_{s})$ and $\boldsymbol{1}(\pi_{i}(X_{i}^{\mathrm{P}})=k)$, that require privatization. We denote the non-private information at time $i$ by
@@
for $B_{s}^{j}\in\mathcal{B}_{i}$ and $k\in[K]$. Specifically, they are privatized as
@@
where $\xi$’s and $\zeta$’s are i.i.d. standard Laplace random variables. The privacy budget $\varepsilon$ is divided into two parts for privacy preservation on $V$’s and $U$’s, respectively.
We remark that all $B_{s}^{j}\in\mathcal{B}_{i}$ receives an update based on $X_{i}^{\mathrm{P}}$ regardless whether $X_{i}^{\mathrm{P}}\in B_{s}^{j}$ or not. Otherwise, bin $B_{s}^{j}$ not receiving an update reveals $X_{i}^{\mathrm{P}}\notin B_{s}^{j}$, which is a privacy leakage. The final estimator is therefore
@@
which satisfies the $\varepsilon$-LDP constraint, as demonstrated in Proposition LABEL:prop:privacyguarantee of the supplement.
@@ 2.3.4 Eliminating Arms
The proposed policy in (8) uniformly pulls all active arms in $A_{s}^{j}$, which implies that we need to exclude arms with large regret from $A_{s}^{j}$. To achieve this, we dynamically rule out arms that are deemed suboptimal in each bin. By a suboptimal arm in a given bin, we mean an arm whose reward function is lower than that of another arm for all $x$ in the bin. Although this is an unobservable population property, it can be inferred using a sufficient condition provided in the following proposition. This proposition establishes a bound between the private estimator (13) and its population counterpart $\mathbb{E}_{Y|X,\pi}\left[\widehat{f}^{\mathrm{P},t}_{k}(x)\right]$, defined as
@@
This result will guide the choice of confidence bound in our arm-elimination procedure.
@@
Let $t_{s}^{j}=\sum_{i=1}^{t}a_{i,s}^{j}$ be defined as in (9). With probability at least $1-n_{\mathrm{P}}^{-2}$, we have for all $t\in[n_{\mathrm{P}}]$ satisfying $t_{s}^{j}\geq\log^{2}(n_{\mathrm{P}})$, it holds that,
@@
for all $k\in A_{s}^{j},B_{s}^{j}\in\mathcal{B}_{t}$, and $x\in B^{j}_{s}\in\mathcal{B}_{t}$, where $C_{n_{\mathrm{P}}}=c\log(n_{\mathrm{P}})$ with a known absolute constant $c$.
The proof of Proposition 2.7 can be found in Section LABEL:sec:error-analysis of the supplement, where we also specify the exact expression of $C_{n_{\mathrm{P}}}$. We remark that Proposition 2.7 gives the very first confidence bound result for LDP nonparametric contextual bandits. In the numerator of (14), the two terms correspond to the private and non-private bounds, respectively. The private term $\varepsilon^{-2}t_{s}^{j}$ arises from the sum of Laplace random variables. As for the non-private term, a more natural form is $\sum_{i=1}^{t}{U}_{i,k,s}^{\mathrm{P},j}a^{j}_{i,s}$, which, however, is unobservable due to the LDP constraints. Since our algorithm requires an accessible realization of $r_{k,s}^{\mathrm{P},t,j}$, we replace this term with its private counterpart $\sum_{i=1}^{t}\tilde{U}_{i,k,s}^{\mathrm{P},j}a^{j}_{i,s}$.
Note that, unlike standard confidence bounds in the nonparametric bandit literature, which count the number of samples whose covariates lie in a particular bin during a given time period, our construction introduces additional Laplacian randomness to comply with LDP requirements. In Lemma LABEL:lem:orderofsumU, we theoretically show that this substitution does not compromise the effectiveness of the confidence bound, provided that $t_{s}^{j}$ is sufficiently large. To ensure this condition is met in practice, we require a sufficient exploration criterion when conducting arm elimination (see (16) below).
An arm elimination rule can be readily derived from (14). In particular, by the triangle inequality, it holds that $|\tilde{f}_{k}^{\mathrm{P},t}(x)-f_{k}(x)|\leq 2r_{k,s}^{\mathrm{P},t,j}$ for all $x\in B_{s}^{j}$ provided that
$\sup_{x\in B_{s}^{j}}\left|\mathbb{E}_{Y|X,\pi}\left[\widehat{f}^{\mathrm{P},t}_{k}(x)\right]-f_{k}(x)\right|\leq r_{k,s}^{\mathrm{P},t,j}.$ (15)
Here, we refer to $\sup_{x\in B_{s}^{j}}\left|\mathbb{E}_{Y|X,\pi}\left[\widehat{f}^{\mathrm{P},t}_{k}(x)\right]-f_{k}(x)\right|$ as the approximation error of bin $B_{s}^{j}$.
Note that condition (15) can be ensured by the bin refinement procedure introduced in the next subsection. Therefore, we can set $2r_{k,s}^{\mathrm{P},t,j}$ as the radius of confidence bound of $\tilde{f}_{k}^{\mathrm{P},t}(x)$ and we eliminate an arm when its upper confidence bound is smaller than the lower confidence bound of another arm. Formally, we remove arm $k^{*}$ from $A_{s}^{j}$ if there exists $k\in[K]$ such that
@@
where the first condition ensures the bin $B_{s}^{j}$ is sufficiently explored, as is required in Proposition 2.7.
@@ 2.3.5 Refining Bins
We now introduce the bin refinement procedure to ensure the claimed condition in (15) holds, which guarantees that the ability to distinguish sub-optimal arms is not dominated by the approximation error of $B_{s}^{j}$. In particular, utilizing the Lipschitz property of the reward function, we choose
@@
as a surrogate for the approximation error. Note that the approximation error is decreasing with $s$, i.e. the finer bins have smaller errors. In fact, (17) is the diameter of $B_{s}^{j}$ and represents an upper bound on the approximation error up to a constant factor, as shown in Lemma LABEL:lem:approximationerror of the supplement. Thus, if $r_{k,s}^{\mathrm{P},t,j}<\tau_{s}$ (i.e. the confidence bound is sufficiently narrow), it signals insufficient approximation capability of the current bin $B_{s}^{j}$, prompting the refinement of the bin using Algorithm 1.
Input: Budget $\varepsilon$. Total sample $n_{\mathrm{P}}.$
Initialization: $\pi_{1}=\operatorname{Unif}([K])$, $\mathcal{B}_{1}=\{B_{0}^{1}\}=\{[0,1]^{d}\}$, $A_{0}^{1}=[K]$.
for $t\in[n_{\mathrm{P}}]$ do
       User side:
       Receive $\pi_{t}$ from the server. Observe $X_{t}^{\mathrm{P}}$, pull arm $\pi_{t}(X_{t}^{\mathrm{P}})$ and receive $Y_{t}^{\mathrm{P},(\pi_{t}(X_{t}^{\mathrm{P}}))}$.
@@
             for $k\in A_{s}^{j}$ do
                   Compute $\tilde{V}_{t,k,s}^{\mathrm{P},j}$ and $\tilde{U}_{t,k,s}^{\mathrm{P},j}$ as in (12) and send to the server. # privatization
             end for
       end for
       Server side:
       for $B_{s}^{j}\in\mathcal{B}_{t}$ do
             for $k\in A_{s}^{j}$ do
                   Update estimates $\tilde{f}_{k}^{\mathrm{P},t}$ as in (13). # estimating reward functions
                   Update confidence bounds as in (14).
             end for
            Remove $k$ from $A_{s}^{j}$ if (16) holds. # eliminating arms
             if  $r_{k,s}^{\mathrm{P},t,j}<{\tau}_{s}$ for some $k\in A_{s}^{j}$  then
@@
                   $\mathcal{B}_{t}=\mathcal{B}_{t}\cup\{B_{s+1}^{2j-1},B_{s+1}^{2j}\}\setminus B_{s}^{j}$. # refining bins
@@
      Set $\mathcal{B}_{t+1}=\mathcal{B}_{t}$, update $\pi_{t+1}$ by (8) and send to the next user.
end for
Algorithm 2 The nonparametric MAB algorithm under LDP
2.3.6 Summary and discussions
Putting things together, Algorithm 2 summarizes the detailed procedure of our proposed algorithm.